OpenStudy (anonymous):

prove that |a+b|<=|a|+|b| @satellite73 @amistre64 @ganeshie8 @myininaya @experimentX @phi @jdoe0001 @thomaster @whpalmer4 just give me idea how to start plz ! anyone !

OpenStudy (anonymous):

square it

OpenStudy (anonymous):

in other words \[|a+b|\leq |a|+|b|\] is can be proved showing that \[(|a+b|)^2\leq (|a|+|b|)^2\]

OpenStudy (anonymous):

so i would get finally (|a+b|)^2 >0 always =a^2+b^2+2ab

OpenStudy (anonymous):

so \[(a+b)^2=a^2+2ab+b^2=|a|^2+2ab+|b|^2\] \[\leq |a|^2+2|a||b|+|b|^2=(|a|+|b|)^2\] as desired

OpenStudy (anonymous):

there is another simple way use the inequalities \[-|a|\leq a\leq |a|\\ -|b|\leq b\leq |b|\] and them up and you get \[-(|a|+|b|)\leq a+b\leq |a|+|b|\] which is identical so saying \[|a+b|\leq |a|+|b|\]

OpenStudy (anonymous):

ok thats make sense thank you !